OpenStudy (anonymous):

Solve for y, then find the value of y when given x = –2. 6x = 7 – 4y

OpenStudy (anonymous):

subtract 7 from both sides, then divide both sides by -4

OpenStudy (anonymous):

so 6x-7=-1 4y-7=-3

OpenStudy (anonymous):

well, follow the step i mentioned above: you are give: 6x=7-4y First subtract 7 from both sides: 6x-7=-4y Now divde both sides by -4 : (6x-7)/(-4)=y

OpenStudy (anonymous):

Now simple replace x with 2

OpenStudy (anonymous):

ok